Let f(x,y)=x^{^{2}}-xy and D be the square given by f(x,y)=x^{^{2}}-xy \left | x \right | \leq 1, \left | y \right |\leq 1. Find the absolute minimum of f(x,y) on D.
Answer #1

We have; + (x, y) = 2²-ny=n(n-y) . Critical points: fn-ty- In = 2n-1 fy= -2 Critical poent in co,o) For absolute value checkiPlease do like.